b1az3:

Divide the following polynomials and then and then complete the quotient. Write your answer in order of decreasing powers of x. (10x^(6) +20x^(4) - 15x^(2)) -: 15x^(2) = ?

jhonyy9:

\[(10x^6 +20x^4 -15x^2 ) : 15x^2 = \] 10x^6 +20x^4 -15x^2 : 15x^2 = 5x^2(2x^4 +4x^2 -3) : 5x^2(3) \[\frac{ 5x^2(2x^4 +4x^2 -3) }{ 5x^2(3) }= ?\] (2x^4 +4x^2 -3) รท 3 = (2/3)x^4 +(4/3)x^2 -1 2x^4 0 0 4x^2 4x^2 0 0 -3 -3 0 hope that s help you

jhonyy9:

but attention pls you need write your answer in order of decreasing of x power